### v1

@@ Abstract
Spatial, kinematic and orbital properties, along with ages and chemical compositions of the thin disc, thick disc and various stellar substructures present in the halo, are studied based on data from the LAMOST and Gaia surveys. The star formation in the Galactic thin and thick disc, with peak metallicities of $-0.20$ and $-0.45$ dex, is found to have peaked about 5.5 and 12.5 Gyr ago, respectively. The thin disc is also found to have experienced an initial star formation burst about 12.5 Gyr ago. The pro-grade population Splash and hot-disc (HD), with peak metallicity of about $-0.60$ and $-0.43$, are found to be about 13.03 and 12.21 Gyr old, respectively, with peak eccentricity of 0.70 and 0.35, are understood to be of in situ origin. The Gaia-Enceladus/Sausage (GE/S), Thamnos and Sequoia, with peak metallicity of about $-1.31$, $-1.36$ and $-1.56$, are found to be about 11.66, 12.89 and 12.18 Gyr, respectively, and are understood to be remnants of dwarf galaxies merged with the Milky Way. The HD, Splash, and Thamnos are found to have experienced chemical evolution similar to the thick disc while GE/S, Sequoia, and Helmi stream are found to have experienced distinct chemical enrichment of iron and $\alpha$-elements.
@@ 2.2 Ages and masses of the sample stars
Figure 1: Colour-magnitude diagram (CDM) and metallicity function for the selected sample of halo stars. A bi-modal distribution is readily visible in the CMD and metallicity function. In the CMD, for ease in identifying the regions where most of the stars are located, we have shown the number density distributions on a logarithmic scale with the ‘cube-helix’ colour schemes in which the dark black colour represents the highest number density while the light grey colour represents the lowest number density.
@@ 3.2 Sample of thin and thick disc stars
Figure 2: Energy-$L_{Z}$ distribution for the kinetically selected sample of halo, thin disc and thick disc stars. In each of the panels, for ease in identifying the regions where most of the sample stars are located, we have shown the number density distributions on a logarithmic scale with the inverted ‘cube-helix’ colour schemes in which the light grey colour represents the highest number density while the dark black colour represents the lowest number density.
@@ 6 Metallicity, age and orbital properties of the Galactic disc’s and halo’s stellar populations
Table 1: Statistical parameters (mean, median, mode, skewness and kurtosis) for distribution functions for all the selected components from the Galactic disc and halo shown in Figure 7 and 8. Here, Mode values are calculated using the well-known empirical relationship $\rm Mode=3\times Median-2\times Mean$. The table also lists the population standard deviation (SD) and the median of absolute deviations from the median (MAD = $median(abs(x-median(x)$) for each of the stellar populations and provides a robust measure of statistical dispersion for each of the stellar populations.
@@
Count Mean SD Median MAD Mode Skew Kurtosis
(dex) (dex) (dex) (dex) (dex)
@@
Count Mean SD Median MAD Mode Skew Kurtosis
(Gyr) (Gyr) (Gyr) (Gyr) (Gyr)
@@
Count Mean SD Median MAD Mode Skew Kurtosis
(kpc) (kpc) (kpc) (kpc) (kpc)
@@ 6.6 Helmi stream (H99)
Figure 11: [$\alpha$/Fe] versus [Fe/H] distributions for the Galactic disc and halo components. In all the panels (as mentioned in the legends), density contours tracing the densest regions of the thin and thick disc are also shown in red and blue colour, respectively. For the substructures in the halo, linear fits to the distributions are also shown and fit parameters are provided in the legends of the corresponding panels.
@@ 7 Chemical evolution of the thin disc, thick disc, and halo’s stellar populations
To find clues about the origins and evolutions of the Galactic disc’s and halo’s populations, we search for the differences in their abundance ratios [X/Fe]. In this work, we restrict our analysis to [$\alpha$/Fe]. A run for [$\alpha$/Fe] as a function of [Fe/H] for all the selected samples of the thin disc, thick disc and components of the Galactic halo is shown in Figure 11. For ease in identifying the regions where most of the stars are located, we have shown the number density distributions on a logarithmic scale with the ‘cube-helix’ colour schemes (as shown in the colour bar). The number density distribution is shown for all populations except for the Thamnos and Sequoia due to their relatively smaller samples. Also, in all the panels (as mentioned in the legends), density contours tracing the densest regions of the thin and thick disc are also shown in red and blue colour, respectively. Additionally, for the substructures in the halo, linear fits to the distributions are also shown, and fit parameters are provided in the legends of the corresponding panels.
@@ 8 Discussion and concluding remarks
All the selected halo sub-structures, except the HD population, are relatively old, with some of their member stars as old as the Universe itself (i.e. about 14 Gyr old). The Splash has a peak age of 13.03 Gyr and is about 1.5 Gyr older than the GE/S, whose peak and median ages are 11.66 and 11.26 Gyr, respectively. Also, most of the Splash and GE/S stars are older than about 8 Gyr. These age estimates are in agreement with many previous studies like Schuster et al. (2012); Hawkins et al. (2014); Ge et al. (2016); however, they are in contrast to a recent study by Gallart et al. (2019) which suggests that both GE/S and Splash (i.e., blue and red colour sequence, respectively) have the same peak age of 13.4 Gyr. For the Splash, our age estimate is in agreement with Gallart et al. (2019)’s age estimate. However, for the GE/S, our age estimate is about 2 Gyr lower than the age estimate provided by Gallart et al. (2019). The disagreement between our age estimates (which are in agreement with many previous studies) and Gallart et al. (2019) appears to be related to the segregation of these two halo’s populations in Gallart et al. (2019). As discussed above and in Section 5, the Galactic halo is home to many substructures, of which the Splash and GE/S are the two major ones. The selection of Splash and GE/S samples purely based on the distribution in the CMD is expected to result in large contamination due to overlap in metallicity and the presence of other small substructures in the halo. Our selection of halo’s sub-structures is based on stars’ kinematics and is independent of their metallicity and distribution in the CMD.
@@
Unlike the thin disc, the thick disc has a more complex distribution of [$\alpha$/Fe] abundances (Figure 11). In the [Fe/H] range from about --2.5 to --0.5 dex, the number of stars with larger [$\alpha$/Fe] at a given [Fe/H] increases with an increase in [Fe/H], suggesting the dominating contribution of $\alpha$-process elements from the SNII. Also, in this metallicity range, the total spread in [$\alpha$/Fe] at a given [Fe/H] is also increasing with an increase in [Fe/H], suggesting the $\alpha$-elements contribution from SNII was not uniform throughout the Galaxy, and the gas was poorly mixed during the formation of the thick disc. Additionally, the presence of stars with lower [$\alpha$/Fe] further confirms the simultaneous contribution of iron from SNIa in some cases. For [Fe/H] $>$ --0.5 dex, the decreasing trend in [$\alpha$/Fe] with respect to an increase in [Fe/H] gives rise to the well-known knee-like distribution for the thick disc and suggests the dominating contribution of iron from SNIa at the later stage of thick disc formation. These trends are in agreement with previous studies, for example, Edvardsson et al. (1993); Reddy et al. (2006). The thin disc is understood to have formed during the merger of a gas-rich system with the progenitor of the Milky Way about 10 Gyr ago (e.g., see Haywood et al., 2018; Hayes et al., 2018; Helmi et al., 2018; Gallart et al., 2019; Koppelman et al., 2019; Belokurov et al., 2020, and references therein). The retrograde component HD and Splash, having [$\alpha$/Fe] abundances similar to the thick disc (except lack of $\alpha$-poor stars), are also believed to have originated during this or a similar merger event (Belokurov et al., 2020). The GE/S, the largest accreted component in the Galactic halo, is known to be the remnant of the last major merger in the Milky Way’s history and has a progenitor mass of about one-fourth of the mass of Milky Way’s progenitor. As shown in Figure 11, the decreasing [$\alpha$/Fe] as a function of [Fe/H] for the GE/S suggests that it GE/S has undergone a distinct chemical evolution than the Milky Way’s thick disc and other in situ components like the HD and Splash. The [$\alpha$/Fe] abundance distribution for the Sequoia and H99 are also similar to the GE/S, suggesting the progenitors of both have undergone similar chemical histories. On the other hand, [$\alpha$/Fe] abundance trend of the Thamnos (which is also believed to be a remnant of a dwarf galaxy) is similar to the Milky Way’s in situ components like HD and Splash (and also metal-poor end of thick disc) and suggests that not all the dwarf galaxies which merged with the Milky Way’s progenitors had chemical evolution distinct than the initial chemical evolution of Milky Way progenitor.

### v2

@@ Abstract
Spatial, kinematic, and orbital properties, along with ages and chemical compositions of the thin disc, thick disc, and various stellar substructures in the halo, are studied based on data from the LAMOST and Gaia surveys. The star formation in the Galactic thin and thick disc, with peak metallicities of $-0.20$ and $-0.45$ dex, is found to have peaked about 5.5 and 12.5 Gyr ago, respectively. The thin disc is also found to have experienced an initial star formation burst about 12.5 Gyr ago. The pro-grade population Splash and hot-disc (HD), with peak metallicity of about $-0.60$ and $-0.43$, are found to be about 13.03 and 12.21 Gyr old, respectively, with peak eccentricity of 0.70 and 0.35, are understood to be of in situ origin. The Gaia-Enceladus/Sausage (GE/S), Thamnos and Sequoia, with peak metallicity of about $-1.31$, $-1.36$ and $-1.56$, are found to be about 11.66, 12.89 and 12.18 Gyr old, respectively, and are understood to be remnants of dwarf galaxies merged with the Milky Way. The HD, Splash, and Thamnos are found to have experienced chemical evolution similar to the thick disc while GE/S, Sequoia, and Helmi stream are found to have experienced distinct chemical enrichment of iron and $\alpha$-process elements.
@@ 2.2 Ages and masses of the sample stars
Figure 1: Colour-magnitude diagram (CDM) and metallicity function for the selected sample of halo stars. A bi-modal distribution is readily visible in the CMD and metallicity function. In the CMD, for ease in identifying the regions where most of the stars are located, we have shown the number density distributions on a logarithmic scale with the ‘cubehelix’ colour schemes in which the dark black colour represents the highest number density while the light grey colour represents the lowest number density (Green, 2011).
@@ 3.2 Sample of thin and thick disc stars
Figure 2: Energy-$L_{Z}$ distribution for the kinetically selected sample of halo, thin disc and thick disc stars. In each of the panels, for ease in identifying the regions where most of the sample stars are located, we have shown the number density distributions on a logarithmic scale with the inverted ‘cubehelix’ colour schemes in which the light grey colour represents the highest number density while the dark black colour represents the lowest number density.
@@ 6 Metallicity, age and orbital properties of the Galactic disc’s and halo’s stellar populations
Table 1: Statistical parameters (mean, median, mode, skewness and kurtosis) for distribution functions for all the selected components from the Galactic disc and halo shown in Figure 7 and 8. Here, Mode values are calculated using the well-known empirical relationship $\rm Mode=3\times Median-2\times Mean$. The table also lists the population standard deviation (SD) and the median of absolute deviations from the median (MAD = $median(abs(x-median(x)$) for each of the stellar populations and provides a robust measure of statistical dispersion for each of the stellar populations. The second column, with the header “Count” lists the number of sample stars used to estimate the statistical parameters for the corresponding substructure.
@@
Mean SD Median MAD Mode Skew Kurtosis
Count (dex) (dex) (dex) (dex) (dex)
@@
Mean SD Median MAD Mode Skew Kurtosis
Count (Gyr) (Gyr) (Gyr) (Gyr) (Gyr)
@@
Mean SD Median MAD Mode Skew Kurtosis
Count (kpc) (kpc) (kpc) (kpc) (kpc)
@@ 6.6 Helmi stream (H99)
Figure 11: [$\alpha$/Fe] versus [Fe/H] distributions for the Galactic disc and halo components. In all the panels (as mentioned in the legend), density contours tracing the densest regions of the thin and thick disc are also shown in red and blue colour, respectively. For the substructures in the halo, linear fits to the distributions are also shown and fit parameters are provided in the legend of the corresponding panels.
@@ 7 Chemical evolution of the thin disc, thick disc, and halo’s stellar populations
To find clues about the origins and evolutions of the Galactic disc’s and halo’s populations, we search for the differences in their abundance ratios [X/Fe]. In this work, we restrict our analysis to [$\alpha$/Fe]. A run for [$\alpha$/Fe] as a function of [Fe/H] for all the selected samples of the thin disc, thick disc and components of the Galactic halo is shown in Figure 11. For ease in identifying the regions where most of the stars are located, we have shown the number density distributions on a logarithmic scale with the ‘cubehelix’ colour schemes (as shown in the colour bar). The number density distribution is shown for all populations except for the Thamnos and Sequoia due to their relatively smaller samples. Also, in all the panels (as mentioned in the legends), density contours tracing the densest regions of the thin and thick disc are also shown in red and blue colour, respectively. Additionally, for the substructures in the halo, linear fits to the distributions are also shown, and fit parameters are provided in the legends of the corresponding panels.
@@ 8 Discussion and concluding remarks
All the selected halo sub-structures, except the HD population, are relatively old, with some of their member stars as old as the Universe itself (i.e. about 14 Gyr old). The Splash has a peak age of 13.03 Gyr and is about 1.5 Gyr older than the GE/S, whose peak and median ages are 11.66 and 11.26 Gyr, respectively. Also, most of the Splash and GE/S stars are older than about 8 Gyr. These age estimates are in agreement with many previous studies like Schuster et al. (2012), Hawkins et al. (2014), Ge et al. (2016), however, they are in contrast to a recent study by Gallart et al. (2019) which suggests that both GE/S and Splash (i.e., blue and red colour sequence, respectively) have the same peak age of 13.4 Gyr. For the Splash, our age estimate is in agreement with Gallart et al. (2019)’s age estimate. However, for the GE/S, our age estimate is about 2 Gyr lower than the age estimate provided by Gallart et al. (2019). The disagreement between our age estimates (which are in agreement with many previous studies) and Gallart et al. (2019) appears to be related to the segregation of these two halo’s populations in Gallart et al. (2019). As discussed above and in Section 5, the Galactic halo is home to many substructures, of which the Splash and GE/S are the two major ones. The selection of Splash and GE/S samples purely based on the distribution in the CMD is expected to result in large contamination due to overlap in metallicity and the presence of other small substructures in the halo. Our selection of halo’s sub-structures is based on stars’ kinematics and is independent of their metallicity and distribution in the CMD.
@@
Unlike the thin disc, the thick disc has a more complex distribution of [$\alpha$/Fe] abundances (Figure 11). In the [Fe/H] range from about --2.5 to --0.5 dex, the number of stars with larger [$\alpha$/Fe] at a given [Fe/H] increases with an increase in [Fe/H], suggesting the dominating contribution of $\alpha$-process elements from the SNII. Also, in this metallicity range, the total spread in [$\alpha$/Fe] at a given [Fe/H] is also increasing with an increase in [Fe/H], suggesting the $\alpha$-elements contribution from SNII was not uniform throughout the Galaxy, and the gas was poorly mixed during the formation of the thick disc. Additionally, the presence of stars with lower [$\alpha$/Fe] further confirms the simultaneous contribution of iron from SNIa in some cases. For [Fe/H] $>$ --0.5 dex, the decreasing trend in [$\alpha$/Fe] with respect to an increase in [Fe/H] gives rise to the well-known knee-like distribution for the thick disc and suggests the dominating contribution of iron from SNIa at the later stage of thick disc formation. These trends are in agreement with previous studies, for example, Edvardsson et al. (1993); Reddy et al. (2006). The thin disc is understood to have formed during the merger of a gas-rich system with the progenitor of the Milky Way about 10 Gyr ago (e.g., see Hayes et al., 2018; Haywood et al., 2018; Helmi et al., 2018; Gallart et al., 2019; Koppelman et al., 2019; Belokurov et al., 2020, and references therein). The retrograde component HD and Splash, having [$\alpha$/Fe] abundances similar to the thick disc (except lack of $\alpha$-poor stars), are also believed to have originated during this or a similar merger event (Belokurov et al., 2020). The GE/S, the largest accreted component in the Galactic halo, is known to be the remnant of the last major merger in the Milky Way’s history and has a progenitor mass of about one-fourth of the mass of Milky Way’s progenitor. As shown in Figure 11, the decreasing [$\alpha$/Fe] as a function of [Fe/H] for the GE/S suggests that it GE/S has undergone a distinct chemical evolution than the Milky Way’s thick disc and other in situ components like the HD and Splash. The [$\alpha$/Fe] abundance distribution for the Sequoia and H99 are also similar to the GE/S, suggesting the progenitors of both have undergone similar chemical histories. On the other hand, [$\alpha$/Fe] abundance trend of the Thamnos (which is also believed to be a remnant of a dwarf galaxy) is similar to the Milky Way’s in situ components like HD and Splash (and also metal-poor end of thick disc) and suggests that not all the dwarf galaxies which merged with the Milky Way’s progenitors had chemical evolution distinct than the initial chemical evolution of Milky Way progenitor.